OpenStudy (anonymous):

I'm really really confused with my algebra one credit recovery, can anyone help me? i don't want to just post the questions constantly, but i need help..

OpenStudy (anonymous):

What's your question?

OpenStudy (anonymous):

Which of the following represents "Twice the abscissa increased by three times the ordinate is ten"? 3x + 2y = 10 2y + 3x = 10 2x + 3y = 10

OpenStudy (anonymous):

@Shames

OpenStudy (anonymous):

2x + 3y = 10

OpenStudy (anonymous):

Describe the location of the point having the following coordinates. negative abscissa, zero ordinate between Quadrant II and Quadrant IV Quadrant II between Quadrant I and Quadrant IV between Quadrant II and Quadrant III i think this one is quad 2..? it wont tell me the ones i miss..

OpenStudy (anonymous):

between quadrant II and III. negative abscissa: x < 0 zero ordinate: y = 0

OpenStudy (anonymous):

ah, got it. Describe the location of the point having the following coordinates. nonzero abscissa, negative ordinate Quadrant II or Quadrant IV Quadrant I or Quadrant II Quadrant I or Quadrant IV Quadrant III or Quadrant IV quad 3 or 4?

OpenStudy (anonymous):

Quadrant I positive abscissa, positive ordinate Quadrant II negative abscissa, positive ordinate Quadrant III negative abscissa, negative ordinate Quadrant IV positive abscissa, negative ordinate

OpenStudy (anonymous):

Describe the location of the point having the following coordinates. abscissa and ordinate have the same sign Quadrant II or Quadrant IV Quadrant I or Quadrant III Quadrant I or Quadrant IV Quadrant II or Quadrant III i just flat out don't understand this one.. && thank you so much for helping me.

OpenStudy (anonymous):

I or III

OpenStudy (anonymous):

and no problem :D

OpenStudy (anonymous):

In which quadrant is (1/2, -1.8)? Quadrant I Quadrant II Quadrant III Quadrant IV wouldn't this one be quad 2?

OpenStudy (anonymous):

no, its IV

OpenStudy (anonymous):

In which quadrant is (-1/2, -1.8)? Quadrant I Quadrant II Quadrant III Quadrant IV im so lost, i don't understand math.

OpenStudy (anonymous):

III

OpenStudy (anonymous):

awsome!!!!!!! thank you so so much.

OpenStudy (anonymous):

Np